PLEASE HELP ASAPA 1000 kg car moving at 5.8 m/s is initially traveling north in the positive y direction. After completing a 90° right-hand t
Answer #1

Initial momentum vector (P1) = 1000*5.8 j = 5800j ( j is unit vector pointing towards y i.e. North)

Assuming it has same speed after turning

after turn momentum (P2) = 1000*5.8 i = 5800i ( i is unit vector pointing towards x i.e. East)  

(a) the magnitude of Impulse = |P2 - P1| = |5800i - 5800j | = 5800*sqrt(2) = 8202.43 N-s

(b) after collision momentum (P3) = 0\

the magnitude of Impulse = |P3 - P1| = |0- 5800i | = 5800 N-s

(c) Average force magnitude = |change in momentum|/time = 8202.43/4.3 = 1907.54 N

(d) Average force = change in momentum/time = 5800/0.220 = 26363.63 N

(e) Average force direction = P2 - P1 = 5800i - 5800j

angle = -atan(5800/5800) = -45 degrees with positive X axis

Please Upvote if you like the answer!!